If 7 men can make 7 chairs in 7 days, how many chairs can 40 men make in 40 days?
I knew the answer before but forgot.
Pls give me the explanation

Answer by LinnW(1048)

If 7 men can make 7 chairs in 7 days,
One man can make (1/7) of 7 chairs in 7 days.
so, one man can make one chair in 7 days.
The rate for one man is one chair/ 7 days.
Dividing the numerator and the denominator by 7
we get (1/7) chairs per day.
So a man can make 1/7 chairs per day.
At this rate we want to know how many chairs can
40 men make in 40 days.
40 men * ( rate for one man ) * ( 40 days )
40 * ( 1/7 per day ) * 40 days
40 * 40 * (1/7) = 1600 * (1/7) =
1600/7 = 228 4/7 chairs